# Compression via Pre-trained Transformers: A Study on Byte-Level Multimodal Data

### v1

@@ 4 Methods
We now describe our experimental setup (with additional details, e.g., sweeps, in Appendix A).
@@
We compare to various standard compressors, both general-purpose, i.e., gzip (Deutsch, 1996) and LZMA2 (Pavlov, 2019), and domain-specific, i.e., FLAC (Coalson, 2008) for audio data and PNG (Boutell, 1997) and lossless JPEG 2000 (Skodras et al., 2001) for images. Both gzip and LZMA2 (which is used by the 7zip software) are based on Huffman coding (Huffman, 1952) and the Lempel-Ziv-Welch algorithm (Welch, 1984). We use the default parameters for gzip, LZMA2, and JPEG 2000, compression level 12 for FLAC, and instruct PNG to find the optimal encoder settings. We also compare to the online transformer from Bellard (2021), with the default v3.3 parameters, which is the current state-of-the-art on the Large Text Compression Benchmark (Mahoney, 2006).
@@
We focus on decoder-only transformers (Vaswani et al., 2017) with SwiGLU activations (Shazeer, 2020) and post-layer normalization. Unless otherwise noted, we use $8$ heads, an embedding dimension of $64$, a context size of $4096$ (bytes), and sliding windows without overlap or memory (full details in Section A.3). We train our models with the Adam optimizer (Kingma and Ba, 2015) for $2.5$ million steps with a batch size of $32$, which, for $165$GB of data, roughly corresponds to $2$ epochs. Due to the duality of compression and prediction, we minimize the standard (sequential) $\log$-loss (Eq. 1) during training, which is a maximum-compression objective (see Section 2).
@@
Tokenization is a commonly-used, domain-specific pre-compression step to boost transformers’ performance by increasing their vocabulary size in order to fit more information into their limited context window (Lester et al., 2024), i.e., increased information density at the cost of increased entropy. However, since our goal is to be domain-general, we do not use tokenization and instead feed our models directly with byte streams (we still have to choose how to flatten images and how to sample audio signals, which are minimal domain-specific preprocessing steps).
@@
$\mathrm{compression\leavevmode\nobreak\ ratio}:=\frac{\mathrm{size\leavevmode\nobreak\ of\leavevmode\nobreak\ compressed\leavevmode\nobreak\ data}+\mathrm{size\leavevmode\nobreak\ of\leavevmode\nobreak\ compressor}}{\mathrm{size\leavevmode\nobreak\ of\leavevmode\nobreak\ uncompressed\leavevmode\nobreak\ data}},$ (2)
which accounts for the model size and is equivalent to the “adjusted compression rate” of Delétang et al. (2024). We always evaluate on $1$GB of out-of-distribution data, i.e., $\mathrm{size\leavevmode\nobreak\ of\leavevmode\nobreak\ uncompressed\leavevmode\nobreak\ data}=1\mathrm{GB}$. As Delétang et al. (2024), we compute the size of the compressor by encoding the model weights with float16 (2 bytes per parameter) since this level of quantization does not significantly affect performance (Tao et al., 2022) and is standard for model inference. As a result, our model sizes range from $0.8$MB to $40.3$MB. Note that, similar to Delétang et al. (2024), we do not compress the model parameters, since naive approaches (e.g., compressing them with gzip) do not significantly decrease the model size (only by around $7$%, which corresponds to a decrease in compression ratio of only $0.002821$ for our largest model). However, as a result, the compression ratio we report is technically an upper bound, which could be improved by (losslessly) compressing the parameters (though with limited room for improvement in our regime, even in the best case).
@@
A key point of our investigation is to evaluate how well pre-trained transformers can compress data from different modalities — both if the modality was or was not part of the training data (Fig. 1 visualizes our data collection process). We create three different unimodal training datasets with audio, images, and text data, and four multimodal training sets (Section A.1 describes the datasets in full detail). This yields seven pre-training datasets in total, each consisting of $165$GB of data: (i) $165$GB of audio; (ii) $165$GB of images; (iii) $165$GB of text; (iv) $82.5$GB of audio and $82.5$GB of images; (v) $82.5$GB of audio and $82.5$GB of text; (vi) $82.5$GB of images and $82.5$GB of text; and (vii) $55$GB audio, $55$GB of images, and $55$GB text. By training our models on all seven training data mixtures, we can investigate in-modality and out-of-modality compression ratios. For example, for a model trained on the text dataset, the in-modality compression ratio can be determined by evaluating on text, while audio or image data provide out-of-modality compression ratios.
Out-of-Distribution Evaluation Datasets
To mimic the setting for which standard compression algorithms were developed (and thereby ensure a fair comparison), where the compressor is programmed with only minimal statistical assumptions about the data (with stronger assumptions for domain-specific compressors), we evaluate on unseen, out-of-distribution (OOD) datasets for each modality and not on in-distribution held-out datasets (as commonly done in machine learning). To do so, we create a single OOD dataset of $1$GB for each modality (full details in Section A.2).
@@ 5 Results
In this section, we present our extensive experimental evaluation (additional results in Appendix B). Unless otherwise noted, we report the best results over two hyperparameter sweeps (described in Section A.3): (i) over the model- and dataset sizes, and (ii) over the model- and context sizes.
Figure 2: Small pre-trained transformers can be domain-general compressors (panels correspond to evaluation data mixtures, bars to training data mixtures). On every out-of-distribution evaluation data mixture, our method (i.e., the bars) outperforms standard compression algorithms (all horizontal lines except for ‘Bellard’) and is on par with Bellard’s online adaptive transformers (the dark blue line) — as long as the evaluation modality was included in the training data mixture. For unseen modalities we observe very little cross-modal transfer (which is different from observations made with foundation models Delétang et al. (2024)). Unimodal training leads to models that are good for their respective modality, but multimodal training yields models that perform almost as well as the unimodal models across all their training modalities (despite seeing a lot less data per modality than the unimodal models), i.e., one can trade off a small amount of performance on each individual modality in return for a strong domain-general compressor via multimodal training (gray bar).
Small Transformers Can Be Domain-General Compressors
Figure 2 shows the best compression ratio attained on each of the seven out-of-distribution evaluation datasets when training a model on each of the seven training data mixtures (we report the best-performing model from our two sweeps for each training-evaluation pair). We observe that transformers can achieve state-of-the-art in-modality compression ratios, regardless of the concrete composition of the training mixture, outperforming standard compression algorithms (even domain-specific ones) in all cases where all evaluation modalities are part of the training mixture. In these cases, transformers thus learn the prototypical statistical patterns related to that modality during pre-training. Importantly, by comparing models trained on unimodal vs. multimodal data, we observe that multimodal training only slightly decreases the compression performance compared to the unimodal models on their respective modalities (despite only having half or a third amount of data from that modality). This means that it is possible to trade off a small amount of performance on each individual modality to obtain a very strong domain-general compressor via multimodal training (the gray bar in Fig. 2).
Figure 3: What you see is what you get. Each panel visualizes the compression ratios for one of our modalities when training models on varying dataset mixtures and sizes. Although one can replace a large proportion of a unimodal training dataset with a multimodal training mixture and not incur a significant loss on the original modality, transformers (at our tested model sizes) do not exhibit improved transfer from the out-of-modality data (i.e., the multimodal models are worse than the unimodal ones, even when trained on much more data from that particular modality). The upshot is that the multimodal training data does not hurt much (note the scale of the y-axis), but leads to significantly improved multimodal compression performance as shown in Fig. 2.
@@
While Fig. 2 shows that substituting half or two thirds of the training set with data from other modalities only leads to a small performance loss compared to the unimodally trained models, it is unclear whether simply training on a smaller amount of unimodal data (i.e., decreasing the unimodal training dataset size to, e.g., $82.5$GB and not substituting $82.5$GB with data from another modality) would give the same performance, or whether there is some transfer between modalities (as suggested by Mirchandani et al. (2023)) that compensates for the smaller amount of data per individual modality. To investigate this, we run an ablation where we subdivide each of our seven training sets into $5$ different sizes: $20\%$, $40\%$, $60\%$, $80\%$, and $100\%$ of the respective dataset (uniformly subsampled). We train a series of models (sweeping over their number of layers; see Section A.3) on each dataset mixture and each dataset size, and then evaluate as before. Figure 3 shows that, for our models and datasets, there is little transfer between modalities. For all cases of audio, text, and (less clearly) images, it is better to train on a smaller unimodal dataset to get the best unimodal performance, as opposed to training on a much larger multimodal dataset. For example, training on a pure text dataset of $33$GB ($20\%$ of $165$GB) outperforms training on a dataset consisting of $82.5$GB (i.e., more than twice as much) text and of $82.5$GB images/audio.
Figure 4: Simultaneously scaling training dataset and model size (for unimodal training- and evaluation data). The colors indicate the model size, and the lines correspond to different dataset sizes ($20\%$, $40\%$, $60\%$, $80\%$, and $100\%$). We always train for $2$ epochs, regardless of dataset size, i.e., smaller datasets require fewer FLOPS. As expected, increasing the number of parameters and the dataset size boosts compression (at the cost of increased training FLOPS). Note that our out-of-distribution evaluation makes models more prone to overfitting, as seen, e.g., for our largest models on images, making scaling more complex than traditionally observed LLM scaling laws.
@@
Since there is a non-trivial relationship between model size and dataset size, we perform a scaling analysis on both of these factors (details in Section A.3). Figure 4 shows trends akin to the scaling laws observed for LLMs (Kaplan et al., 2020), which state that better prediction (in our case compression) is only possible by scaling both models and datasets, in a particular way. Note that, different to traditional scaling laws for models trained on internet-scale datasets, the distribution shift in our evaluation makes it easier for the model to overfit to the training distribution. However, as the number of parameters and the training flops of our small models increase, the adjusted compression ratio improves, eventually beating standard compression algorithms. We do observe gradual overfitting on the image dataset for our models trained only on images. However, this phenomenon can be mitigated by including other modalities in the training mixture (see Fig. A1).
@@
Figure 5: Relationship between context- and model size. Given a certain training compute budget (in FLOPS), one can either increase the context size (measured in bytes) or the model size, leading to a non-trivial trade-off. Our results show that this trade-off is highly modality-dependent (also note the different scales on the y-axis, meaning that the magnitude of the effect varies significantly with modality). For text, shorter context sizes and larger models are beneficial (indicating the importance of short-term dependencies for our data and model scale). For images, larger context is generally beneficial, which makes sense, given that a single image consists of $512\cdot 512\cdot 3=786432$ bytes, which far exceeds our models’ contexts, i.e., models with larger context can process larger fractions of an image. Finally, for audio data the relationship is complex with intermediate context length and larger models performing better (though the reverse is true for short context length).
The previous two experiments investigated the impact of training dataset size and model size, which revealed a complex, “scaling law”-like, relationship between the two factors and the overall training budget in FLOPS. In this experiment, we investigate the impact of the length of the context window. Since the context window length has a large impact on the overall FLOPS footprint (attention scales quadratically with the input sequence length), we also vary the size of our models to explore whether there is a sweet spot in terms of training compute budget allocation (details in Section A.3). Fig. 5 shows that the optimal trade-off strongly depends on the data modality. The top performing models for text have a context window less than or equal to $2048$ bytes, indicating that short term dependencies are more important than long ones in this case. For images, the best compromise overall is to choose a larger context window of $8192$, which means decreasing the model size. For audio data, the trade-off is even more complex. Overall these results highlight the difficulty of tuning architectures to achieve best performance across many modalities.
Figure 6: Impact of the sliding window overlap (for unimodal training and evaluation). Having overlapping context windows only leads to marginal performance increases (most significant for images) in our experiments but comes at a huge cost in terms of computational efficiency.
@@
In all experiments so far we used a sliding window without overlap to process the evaluation byte streams, i.e., we completely fill a whole context window, process it, and then slide it forward by the size of the context window to process the next chunk of data. This means that bytes early in the context window are not conditioned on a lot of data (in theory, conditioning on more data should help with prediction and thus compression, which may well be exploitable by transformers’ in-context learning abilities (Genewein et al., 2023; Ge et al., 2024)). However, sliding the context window with more overlap requires more forward passes to process the same amount of data, which significantly increases the computational cost with increasing overlap. With no overlap processing $4096$ bytes with a context window of $4096$ takes a single forward pass. In the most extreme case of maximal overlap it would take $4095$ forward passes, where the context window is moved by a single byte each time (though each prediction could be conditioned on the full $4095$ preceding bytes). In our final experiment, we investigate the effect of different overlaps between context windows. Figure 6 shows that for our data and model sizes, increasing the overlap window (for a context length of 4096) has relatively little effect. The strongest effect is observed for image data, which makes sense given that 4096 bytes only corresponds to a small fraction of an image and there are obvious long-range dependencies between channels of the same image. Beyond an overlap of $2048$ we do not see much benefit of further increasing the overlap window in our experiments.

### v2

@@ 4 Methods
We now describe our experimental setup (cf. Appendix B).
@@
We compare to various standard compressors, both general-purpose, i.e., gzip (Deutsch, 1996) and LZMA2 (Pavlov, 2019), and domain-specific, i.e., FLAC (Coalson, 2008) for audio data and PNG (Boutell, 1997) and lossless JPEG 2000 (Skodras et al., 2001) and JPEG-XL (Alakuijala et al., 2019) for images. Both gzip and LZMA2 (which is used by the 7zip software) are based on Huffman coding (Huffman, 1952) and the Lempel-Ziv-Welch algorithm (Welch, 1984). We use the default parameters for gzip, LZMA2, JPEG 2000, and JPEG-XL, compression level 12 for FLAC, and instruct PNG to find the optimal encoder settings. We also compare to the online transformer from Bellard (2021) with the default v3.3 parameters, which is currently state-of-the-art on the Large Text Compression Benchmark (LTCB) (Mahoney, 2006). We do not compare to Izacard et al. (2020), since it is conceptually very close to Bellard (2021) but worse on LTCB.
@@
We focus on decoder-only transformers (Vaswani et al., 2017) with SwiGLU activations (Shazeer, 2020) and post-layer normalization. Unless otherwise noted, we use $8$ heads, an embedding dimension of $64$, a context size of $4096$ (bytes), and sliding windows without overlap or memory (full details in Section B.3). We always train and evaluate the models with the same context size ($4096$ by default). We use the Adam optimizer (Kingma & Ba, 2015) for $2.5$ million steps with a batch size of $32$, which, for $165$GB of data, roughly corresponds to $2$ epochs. Due to the duality of compression and prediction, we minimize the standard (sequential) $\log$-loss (Eq. 1) during training, which is a maximum-compression objective (see Section 2). The model computes a distribution over the next byte for every input byte, and the arithmetic coder then uses these predictions to losslessly compress the data. Concretely, the arithmetic coder directly uses the model’s predictions over tokens (i.e., the logits) to encode/decode data (see Figure 1 in Delétang et al. (2024) for an overview). Accordingly, no separate head or extraction procedure is necessary. As a result, we can train via standard log-loss minimization to perform next-byte prediction. At inference time, we perform standard autoregressive evaluation (using teacher forcing).
@@
Tokenization is a commonly-used, domain-specific pre-compression step to boost transformers’ performance by increasing their vocabulary size to fit more information into their context window (Lester et al., 2024), i.e., increased information density at the cost of increased entropy. However, since we aim to be domain-general, we do not use tokenization and instead feed byte streams to our models (we still have to choose how to flatten images/sample audio signals, i.e., minimal domain-specific preprocessing).
Figure 2: Small pre-trained transformers are domain-general compressors (panels: evaluation data mixtures, bars: training data mixtures). Our method (bars) outperforms standard compression algorithms (horizontal lines) and is on par with the online adaptive transformers from Bellard (2021) (blue line) — as long as the evaluation modality is in the training mixture. There is very little cross-modal transfer to unseen modalities (unlike foundation models (Delétang et al., 2024)). Unimodal models are good for their respective modality, but multimodal models perform almost as well across all their training modalities (despite seeing much less data per modality than the unimodal models), i.e., one can trade off a small amount of performance on each individual modality to obtain a strong domain-general compressor via multimodal training (gray bar).
@@
$\small\mathrm{compression~{}ratio}:=\frac{\mathrm{|compressed~{}data|}+\mathrm{|compressor|}}{\mathrm{|uncompressed~{}data|}},$ (3)
which accounts for the model size and is equivalent to the “adjusted compression rate” of Delétang et al. (2024). We always evaluate on $1$GB of out-of-distribution data, i.e., $|\mathrm{uncompressed~{}data}|=1\mathrm{GB}$. Like Delétang et al. (2024), we compute the size of the compressor by encoding the model weights with float16 (2 bytes per parameter) since this level of quantization has little impact on performance (Tao et al., 2022) and is standard for model inference. As a result, our model sizes range from $0.8$MB to $40.3$MB. Note that, similar to Delétang et al. (2024), we do not compress the model parameters, since naive approaches (e.g., compressing them with gzip) barely decrease the model size (only by around $7$%, which corresponds to a decrease in compression ratio of only $0.002821$ for our largest model). However, as a result, the compression ratio we report is an upper bound, which could be improved by (losslessly) compressing the parameters (though with limited room for improvement in our regime, even in the best case).
@@
A key point of our investigation is to evaluate how well pre-trained transformers can compress data from different modalities — both if the modality was or was not part of the training data (see Fig. 1). We create three different unimodal training datasets with audio, images, and text, and four multimodal training sets (full details in Section B.1). This yields seven $165$GB pre-training datasets: (i) $165$GB audio; (ii) $165$GB images; (iii) $165$GB text; (iv) $82.5$GB audio and $82.5$GB images; (v) $82.5$GB audio and $82.5$GB text; (vi) $82.5$GB images and $82.5$GB text; and (vii) $55$GB audio, $55$GB images, and $55$GB text. By training models on all seven data mixtures, we can investigate in-modality and out-of-modality compression ratios. For example, for a model trained on text (iii), the in-modality compression ratio is given by evaluating on text, while audio or image data provide out-of-modality compression ratios.
OOD Evaluation Datasets
To mimic the standard compression algorithm setting (and thereby ensure a fair comparison), where the compressor is designed with only minimal statistical assumptions about the data (domain-specific compressors make stronger assumptions), we evaluate on unseen, out-of-distribution (OOD) datasets for each modality and not on in-distribution held-out datasets (as common in machine learning). To do so, we create a single OOD dataset of $1$GB for each modality (details in Section B.2).
Figure 3: What you see is what you get. Each panel visualizes the compression ratios for one of our modalities when training models on varying dataset mixtures and sizes. Although one can replace a large proportion of the unimodal training datasets with data from other modalities without incurring significant losses on the original modality (note the scale of the y-axis), transformers (at our tested model sizes) do not exhibit improved transfer from the out-of-modality data (i.e., the multimodal models are worse than the unimodal ones, even when trained on much more data from that particular modality). Nevertheless, multimodal training data significantly improves multimodal compression performance (as shown in Fig. 2).
@@ 5 Results
We present our extensive experimental evaluation (see also Appendix C). Unless otherwise noted, we report the best results over two hyperparameter sweeps (see Section B.3): (i) model- vs. dataset size, and (ii) model- vs. context size.
Small Transformers Are Domain-General Compressors
Figure 2 shows the best compression ratio attained on each of the seven out-of-distribution evaluation datasets when training a model on each of the seven training data mixtures (we report the best-performing model from our two sweeps for each training-evaluation pair). We observe that pre-trained transformers achieve state-of-the-art in-modality compression ratios, regardless of the concrete composition of the training mixture, outperforming standard compression algorithms (even domain-specific ones) in all cases where all evaluation modalities are part of the training mixture. In these cases, transformers thus learn the modality’s prototypical statistical patterns during pre-training. Importantly, by comparing models trained on unimodal vs. multimodal data, we observe that multimodal training only slightly decreases performance compared to the unimodal models on their respective modalities (despite only having half or a third amount of data from that modality). This means that it is possible to trade off a small amount of performance on each individual modality to obtain a very strong domain-general compressor via multimodal training (the gray bar in Fig. 2).
Figure 4: Scaling training dataset- and model size (for unimodal training and evaluation). Colors indicate the model size; lines correspond to dataset size. We train for $2$ epochs regardless of dataset size (i.e., smaller datasets require fewer FLOPS). Increasing the model- and dataset size boosts compression (at the cost of FLOPS). Our OOD evaluation makes models more prone to overfitting (e.g., our largest image models), making scaling more complex than traditional LLM scaling laws.
@@
While Fig. 2 shows that substituting half or two thirds of the training set with data from other modalities only leads to a small performance loss compared to the unimodally trained models, it is unclear whether simply training on a smaller amount of unimodal data (i.e., decreasing the unimodal training dataset size to, e.g., $82.5$GB, and not substituting $82.5$GB with another modality) would give the same performance, or whether there is some transfer between modalities (as suggested by Mirchandani et al. (2023)) that compensates for the smaller amount of data per individual modality. To investigate this, we run an ablation where we subdivide each of our seven training sets into $5$ different sizes: $20\%$, $40\%$, $60\%$, $80\%$, and $100\%$ of the respective dataset (uniformly subsampled). We train a series of models (sweeping over their number of layers; see Section B.3) on each dataset mixture and each dataset size, and then evaluate as before. Figure 3 shows that, for our models and datasets, there is little transfer between modalities. For all cases of audio, text, and (less clearly) images, it is better to train on a smaller unimodal dataset to get the best unimodal performance, as opposed to training on a much larger multimodal dataset. For example, training on a pure text dataset of $33$GB ($20\%$ of $165$GB) outperforms training on a dataset consisting of $82.5$GB (i.e., more than twice as much) text and of $82.5$GB images/audio.
@@
Since there is a non-trivial relationship between model- and dataset size, we analyze the scaling of these factors (details in Section B.3). Figure 4 shows trends akin to the scaling laws observed for LLMs (Kaplan et al., 2020), which state that better prediction (in our case compression) is only possible by scaling both models and datasets in a particular way. Unlike traditional scaling laws for models trained on internet-scale datasets, the distribution shift in our evaluation makes it easier for the model to overfit to the training distribution (since we always evaluate on OOD data). However, as the number of parameters and the training flops of our small models increase, the adjusted compression ratio improves, eventually beating standard compression algorithms. We do observe gradual overfitting on the image dataset for our models trained only on images. However, this phenomenon can be mitigated by including other modalities in the training mixture (see Fig. A1).
Figure 5: Context- vs. model size. Both context size (measure in bytes) and model sizes affect the training compute budget (in FLOPS), leading to a non-trivial trade-off. Our results show that this trade-off is highly modality-dependent (note the different y-axis scales, i.e., the effect varies significantly with modality). For text, shorter context sizes and larger models are beneficial (short-term dependencies are most important). For images, larger context is generally beneficial, given that a single image consists of $512\cdot 512\cdot 3=786432$ bytes, far exceeding our models’ contexts, i.e., models with larger context can process larger fractions of an image at once. For audio, the relationship is complex with intermediate context length and larger models performing better (the reverse is true for short contexts).
Figure 6: Impact of the sliding window overlap (for unimodal training/evaluation). Increasing the overlap between context windows only marginally improves the performance (most significantly for images) but comes at a huge computational cost.
@@
The previous two experiments investigated the impact of training dataset- and model size, revealing a complex, “scaling law”-like, relationship between the two and the overall FLOPS training budget. We now investigate the impact of the context window length. Since the context window length has a large impact on the overall FLOPS footprint (attention scales quadratically with input sequence length), we also vary the model size to explore whether there is a sweet spot in terms of training compute budget allocation (details in Section B.3). Fig. 5 shows that the optimal trade-off strongly depends on the data modality. The best models for text have a context window $\leq 2048$ bytes, indicating that short term dependencies are more important than long ones in this case. For images, the best compromise overall is to choose a larger context window of $8192$, which means decreasing the model size. For audio data, the trade-off is even more complex. Overall these results highlight the difficulty of tuning architectures to achieve best performance across many modalities.
@@
So far we used a sliding window without overlap to process the evaluation data. Consequently, bytes early in the context window are not conditioned on a lot of data (conditioning on more data should help with prediction and thus compression, according to transformers’ in-context learning abilities (Brown et al., 2020; Genewein et al., 2023; Ge et al., 2024)). Sliding the context window with more overlap requires more forward passes to process the same amount of data, significantly increasing the computational cost. Figure 6 shows that increasing the overlap window (for a context length of $4096$) has relatively little effect. The strongest effect is observed for image data, since $4096$ bytes only captures a small fraction of an image and there are obvious long-range dependencies between image channels.
Evaluation Dataset Size
Figure 7: Compression ratio vs. evaluation dataset size. The numerator of the compression ratio consists of the size of the compressed data and the size of the compressor (Eq. 3). For standard compressors (e.g., gzip), the size of the compressor (a few thousand lines of code) is negligible given sufficient evaluation data (i.e., the compression ratio is unaffected by the evaluation dataset size). For neural compressors trained offline (i.e., where the size of the compressor is dominated by the model parameters), the compression ratio improves with increasing data since the model size influence decreases. If the model size $\geq$ evaluation dataset size (e.g., $500$M parameters and $1$GB of data), one cannot achieve a compression ratio $<1$.
Figure 7 shows the relationship between the compression ratio and the evaluation dataset size for all three modalities and our best-performing model (as determined on $1$GB of OOD data in Table A2). For offline (i.e., pre-) trained neural compressors, the model parameters are factored into the compression ratio, which means that their compression performance will improve with increasing evaluation data (as long as the model generalizes well to the additional data). In contrast, the size of standard compressors is negligible compared to the amount of evaluation data, which means that their compression ratios are largely unaffected by the evaluation dataset size. FLAC cannot losslessly compress more than $4.2$GB of data.